Type or paste question here

1. EX.12.01.ALGO

eBook

Show Me How

Average Rate of Return

The following data are accumulated by Watershed Inc. in evaluating two competing capital investment proposals:

Project A

Project Z

Amount of investment

$88,000

$92,000

Useful life

4 years

9 years

Estimated residual value

0

0

Estimated total income over the useful life

$14,080

$53,820

Determine the expected average rate of return for each project. Round your answers to one decimal place.

Project A

%

Project Z

%

2. EX.12.03.ALGO

eBook

Show Me How

Average Rate of Return—New Product

Hana Inc. is considering an investment in new equipment that will be used to manufacture a smartphone. The phone is expected to generate additional annual sales of 3,900 units at $214 per unit. The equipment has a cost of $326,400, residual value of $24,600, and an 8-year life. The equipment can only be used to manufacture the phone. The cost to manufacture the phone follows:

Cost per unit:

Direct labor

$36.00

Direct materials

140.00

Factory overhead (including depreciation)

24.05

Total cost per unit

$200.05

Determine the average rate of return on the equipment. If required, round to the nearest whole percent.
%

3. EX.12.04.ALGO

eBook

Show Me How

Determine Cash Flows

Natural Foods Inc. is planning to invest in new manufacturing equipment to make a new garden tool. The new garden tool is expected to generate additional annual sales of 7,900 units at $32 each. The new manufacturing equipment will cost $102,700 and is expected to have a 10-year life and a $7,900 residual value. Selling expenses related to the new product are expected to be 5% of sales revenue. The cost to manufacture the product includes the following on a per-unit basis:

Direct labor

$5.40

Direct materials

17.90

Fixed factory overhead-depreciation

1.20

Variable factory overhead

2.70

Total

$27.20

Determine the net cash flows for the first year of the project, Years 2–9, and for the last year of the project. Use the minus sign to indicate cash outflows. Do not round your intermediate calculations but, if required, round your final answers to the nearest dollar.

Natural Foods Inc.

Net Cash Flows

Year 1

Years 2-9

Last Year

Initial investment

$

Operating cash flows:

Annual revenues

$

$

$

Selling expenses

Cost to manufacture

Net operating cash flows

$

$

$

Total for Year 1

$

Total for Years 2–9 (operating cash flow)

$

Residual value

Total for last year

$
Answer #1

Average rate of return = Annual net income/Average investment

Project A

= (14080/4) / (88,000/2)

8%

Project Z

= (53820/9)/(92,000/2)

13%

Answer 2.

Average rate of return = Annual net income/Average investment

Average rate of return = [ [214-200.05] * 3900] / [(326400+24600) / 2] = 54405/175500 = 31%

note: as per HomeworkLib policy, first question is answered.